For the vibrating system shown in Fig. 3, a mass of 5 kg is placed on mass m at t = 0 and the system is at rest initially (at t = 0). Given that m = 20 kg, k = 600 N/m, and c = 60 Ns/m. Plot the response curve x(t) versus t using MATLAB.

3 kg m Fig. 3
Answer #1

F(t): 5x9.81 = 49.05N M = m+ 3+5. a - 28 kg x boir m+8 = 20+8 M = 28 kg ↑ book The system is rest. .: 210) -0, pc10 -0. DAle0.14 0.12 0.1 0.08 x(t) in meter 0.06 0.04 0.02 0 0 1 2 3 4 6 7 8 9 10 5 time in sec clc;clear all;close all; T=0:0.001:10;%t